q1. Find the area under the Standard Normal Curve in-between z = -0.63 and z = 0.47

Use the Normal table and give answer using 4 decimal places.

q2. Find the area under the Standard Normal Curve in-between z = -1.19 and z = 0.36

Use the Normal table and give answer using 4 decimal places.

q3. Find P(z < 1.72)

Use the Normal table and give answer using 4 decimal places.

q4. Find P(z > -2.62)

Use the Normal table and give answer using 4 decimal places.
Answer #1

Solution,

Using standard normal table,  

1) P( -0.63 < Z < 0.47)

= P( Z < 0.47) - P( Z < -0.63)

= 0.6808 - 0.2643

= 0.4165

2) P( -1.19 < Z < 0.36)

= P( Z < 0.36) - P( Z < -1.19)

= 0.6406 - 0.1170

= 0.5236

3) P( Z < 1.72)

= 0.9573

4) P(Z > -2.62)

= 1 - P( Z < -2.62)

= 1 - 0.0044

= 0.9956